# Top Quark Mass Extractions from Energy Correlators: A Feasibility Study

### v1

@@ 4.2.1 Parton Distribution Functions
In figure -1011, we show the results of the PDF variations, which were found to have no significant impact.
@@ 4.2.2 Initial and Final State Radiation
These variations are shown in figure -1010. Both variations were found to be not significant, however, notably, the effects were more prominent for FSR ($\approx 100$ MeV). This is expected since varying FSR modified the final state perturbative description of the top decay, to which our observable is sensitive by design.
@@ 4.2.3 Matching Corrections to Stable Top Production
We now focus on the impact of next-to-leading order (NLO) matching to the hard-process production of stable top quarks. Physically, this matching procedure modifies only the amount of gluonic radiation from the hard process before the top is produced and, therefore, the overall boost applied to the produced top quark. However, the boost of the top quark and the boost of the $W$ boson are strongly correlated, which is why the jet $p_{T}$ dependence cancels in the ratio of their peaks. Therefore, the extracted top mass is expected to be very stable under variation of the NLO matching. This is confirmed in our analyses. Results are shown in figure -1009. Including NLO corrections to the hard scattering has no significant impact. Again, this focuses the theoretical attention on improving the description of the top quark decay.
Figure -1009: Impact of NLO corrections to the stable top quark production process.
@@ 4.2.4 Jet Radius
Figure -1008: Effects of the variation of the top jet radius. The grey hashed region illustrates the region, $R\lesssim\sqrt{2\pi}\,m_{t}/p_{T,\mathrm{jet}}$, where jet radius effects are non-negligible.
In figure -1008, we show variations for different values of the top jet radius. For very small jet radii such as $R=0.8$, the cone size is insufficient for lower $p_{T,\mathrm{jet}}$ bins. This can be seen as a significant sensitivity of the ratio to the underlying event contamination and a mild $p_{T,\mathrm{jet}}$ dependence. However, for larger jet radii $R>1$, we find excellent robustness against hadronization and MPI effects, as well as a flat dependence on $p_{T,\mathrm{jet}}$. For $R=1.2$ and $R=1.5$, we find a shift from hadronization and MPI, averaged over all $p_{T,\mathrm{jet}}$ values shown, to be
@@
where the uncertainty is statistical. The common trend displayed in figure -1008 as well as the compatibility of the shifts for the different values of $R$ together with the agreement of the results for different values of the perturbative coefficient $C$ demonstrates that the dependence of the ratio observable on $R$ can be captured entirely in perturbation theory.
@@ 4.2.5 Perturbative Top Decay and Showering
The sensitivities to the extracted top mass are shown in figure -1007 for differing recoil schemes. A 500500500~{}500MeV shift is observed, which, as explained above, is of purely perturbative origin. Table 1 shows the change in the parameter $C$ with different perturbative parton shower models. Notably, the three dipole showers, which each handle the NLO top decay phase-space in an equivalent fashion, are consistent within ±$\pm 0.5\%$. In contrast, the Herwig angular ordered parton shower does not fill the complete NLO phase space and computes a value of $C$ larger by about $2\%$. This is consistent with the expectation from factorization, which is that underestimating the NLO contribution will lead to an overall smaller characteristic decay angle for the top quark.
Figure -1007: Impact of the choice of recoil scheme in the top quark decay.
Figure -1006: Impact of hadronization corrections in different parton shower generators.
@@ 4.2.6 Hadronization
In figure -1006, we show the impact of hadronization corrections for different parton shower generators and for various values of $p_{T,\text{jet}}$. The observable has a minimal sensitivity of order 100100100~{}100MeV to hadronization corrections. This suggests that nonperturbative power corrections are under good control and confirms the results presented in Ref. [27].
@@ 4.2.7 $b$-fragmentation
Figure -1005: Impact of variations in the parameterization of the $b$-fragmentation function.
@@
Sensitivity to the b-quark fragmentation in top quark mass extraction studies is often obtained by varying the Pythia parameter StringZ:rFactB. From measurements at LEP, this parameter is tuned to a best-fit value of StringZ:rFactB = 1.056. In studies at CMS, it is varied between StringZ:rFactB = 0.855 and StringZ:rFactB = 1.252 to test the impact. The results of this variation are shown in figure -1005, where negligible impact is found. This illustrates insensitivity to the modelling of the b-quark fragmentation process.

### v2

@@ 4.2.1 Parton Distribution Functions
In figure -950, we show the results of the PDF variations, which were found to have no significant impact.
@@ 4.2.2 Initial and Final State Radiation
These variations are shown in figure -949. Both variations were found to be not significant, however, notably, the effects were more prominent for FSR ($\approx 100$ MeV). This is expected since varying FSR modified the final state perturbative description of the top decay, to which our observable is sensitive by design.
@@ 4.2.3 Matching Corrections to Stable Top Production
We now focus on the impact of next-to-leading order (NLO) matching to the hard-process production of stable top quarks. Physically, this matching procedure modifies only the amount of gluonic radiation from the hard process before the top is produced and, therefore, the overall boost applied to the produced top quark. However, the boost of the top quark and the boost of the $W$ boson are strongly correlated, which is why the jet $p_{T}$ dependence cancels in the ratio of their peaks. Therefore, the extracted top mass is expected to be very stable under variation of the NLO matching. This is confirmed in our analyses. Results are shown in figure -948. Including NLO corrections to the hard scattering has no significant impact. Again, this focuses the theoretical attention on improving the description of the top quark decay.
Figure -948: Impact of NLO corrections to the stable top quark production process.
@@ 4.2.4 Jet Radius
Figure -947: Effects of the variation of the top jet radius. The grey hashed region illustrates the region, $R\lesssim\sqrt{2\pi}\,m_{t}/p_{T,\mathrm{jet}}$, where jet radius effects are non-negligible.
In figure -947, we show variations for different values of the top jet radius. For very small jet radii such as $R=0.8$, the cone size is insufficient for lower $p_{T,\mathrm{jet}}$ bins. This can be seen as a significant sensitivity of the ratio to the underlying event contamination and a mild $p_{T,\mathrm{jet}}$ dependence. However, for larger jet radii $R>1$, we find excellent robustness against hadronization and MPI effects, as well as a flat dependence on $p_{T,\mathrm{jet}}$. For $R=1.2$ and $R=1.5$, we find a shift from hadronization and MPI, averaged over all $p_{T,\mathrm{jet}}$ values shown, to be
@@
where the uncertainty is statistical. The common trend displayed in figure -947 as well as the compatibility of the shifts for the different values of $R$ together with the agreement of the results for different values of the perturbative coefficient $C$ demonstrates that the dependence of the ratio observable on $R$ can be captured entirely in perturbation theory.
@@ 4.2.5 Perturbative Top Decay and Showering
The sensitivities to the extracted top mass are shown in figure -946 for differing recoil schemes. A 500500500~{}500MeV shift is observed, which, as explained above, is of purely perturbative origin. Table 1 shows the change in the parameter $C$ with different perturbative parton shower models. Notably, the three dipole showers, which each handle the NLO top decay phase-space in an equivalent fashion, are consistent within ±$\pm 0.5\%$. In contrast, the Herwig angular ordered parton shower does not fill the complete NLO phase space and computes a value of $C$ larger by about $2\%$. This is consistent with the expectation from factorization, which is that underestimating the NLO contribution will lead to an overall smaller characteristic decay angle for the top quark.
Figure -946: Impact of the choice of recoil scheme in the top quark decay.
Figure -945: Impact of hadronization corrections in different parton shower generators.
@@ 4.2.6 Hadronization
In figure -945, we show the impact of hadronization corrections for different parton shower generators and for various values of $p_{T,\text{jet}}$. The observable has a minimal sensitivity of order 100100100~{}100MeV to hadronization corrections. This suggests that nonperturbative power corrections are under good control and confirms the results presented in Ref. [27].
@@ 4.2.7 $b$-fragmentation
Figure -944: Impact of variations in the parameterization of the $b$-fragmentation function.
@@
Sensitivity to the b-quark fragmentation in top quark mass extraction studies is often obtained by varying the Pythia parameter StringZ:rFactB. From measurements at LEP, this parameter is tuned to a best-fit value of StringZ:rFactB = 1.056. In studies at CMS, it is varied between StringZ:rFactB = 0.855 and StringZ:rFactB = 1.252 to test the impact. The results of this variation are shown in figure -944, where negligible impact is found. This illustrates insensitivity to the modelling of the b-quark fragmentation process.